A tuning fork vibrates 200 times per second and sends out a compressional wave



that travels with a speed of 340 m/s. Find (a) the period and (b) the wavelength.

1
Expert's answer
2022-03-01T11:12:20-0500

Given:

"f=200\:\rm Hz"

"v=340\:\rm m/s"

(a) the period

"T=\frac{1}{f}=\frac{1}{200}=0.005\:\rm s"

(b) the wavelength

"\lambda=\frac{v}{f}=\frac{340}{200}=1.7\:\rm m"